OpenStudy (anonymous):

Solving to step Equations: Chuck is making a sand box for a park. He knows the width of the sandbox is 3 1/2 feet long, and the perimeter is 16.5 feet. What us the length of the sandbox?( please use variables)

OpenStudy (anonymous):

sandbox= square or rectangle. so 3.5+3.5= 7 16.5-7=9.5 9.5/2=4.75 or 4 3/4 feet.

OpenStudy (anonymous):

You can just write it out word by word: length times 2 plus width times 2 equals perimeter 2l+2p = perimeter then just plug in your numbers for length and perimeter and use algebra to solve

OpenStudy (radar):

You want variables. Then let l= length of sandbox. (variable l) The width is given w=3.5 ft. You also know the perimeter is P=2l + 2w Plug in known and variable. 16.5=2(l)+2(3 1/2) 16.5=2(l) + 7 2(l)=16.5-7=9.5 ft l=9.5/2=4.75 length is 4.75 Now check. 2(4.75)+2(3.5)=16.5 ft. 9.5+7+16.5 16.5=16.5 It checks out.

OpenStudy (radar):

Did you follow that with understanding?